OpenStudy (anonymous):

One of the endpoints of the base of a triangle is at P(4, 5). The midpoint of the base is at (1, 1). Erik performed the steps shown to find the coordinates of the other endpoint of the base. Step 1. x1 = 4, y1 = 5, x = 1, y = 1 Step 2. Step 3. Step 4. 2 = 4 + x2, 2 = 5 + y2 Step 5. 0 = 4 + x2 , 0 = 5 + y2 Step 6. x2 = -4, y2 = -5 In which step did Erik first make an error? Answer Step 2, because he used the formula to find the length of a line segment. Step 3, because he substituted the values incorrectly in the formula Step 4, because he multiplied both the equat

OpenStudy (anonymous):

does step 4 pr 5 look wronf i cant figure out how t owrite out step 2 and 3

OpenStudy (amistre64):

how would you determine the midpoint? match that to the ficticious persons method and see what fits

OpenStudy (amistre64):

or end points or whatever the questions asking ....

OpenStudy (anonymous):

step two is the midpoint between two point formula and step three is the same but puting in the numbers

OpenStudy (amistre64):

Midpoint - endpoint given is the distance and direction to the end point: (M-P) add that result to the midpoint to go the same distance and direction further: M+(M-P) = 2M-P for the other end point

OpenStudy (anonymous):

so step 4 would be wrong?

OpenStudy (amistre64):

2M = (2 , 2) - P = (-4,-5) ------------- (-2,-3) is what I get

OpenStudy (anonymous):

okay but i have to find out what step is wrong thats what i dont get

OpenStudy (amistre64):

i dont do the ficticious persons steps, so I really cant say wear they messed up at .... and it would help to know what step 2 and 3 actually were

OpenStudy (amistre64):

maybe a formula?

OpenStudy (amistre64):

\[\left(\frac{Mx-Px}{2}, \frac{My-Py}{2}\right)\]

OpenStudy (amistre64):

that doesnt quite jive ....

OpenStudy (amistre64):

\[(Mx,My)=\left(\frac{Px+Ex}{2}, \frac{Py+Ey}{2}\right)\] that might seem better

OpenStudy (amistre64):

M for midpt, P for given Point, and E for other endpt

OpenStudy (anonymous):